AS Given the values of AH and T below, determine Ixn rxn ASuniv
Part A 133 kJ, AS =250 J/K,T = 294 K ΔΗ гxn Express your answer using three significant figures. spontaneous or nonspontaneous

Part B н. 133 kJ, AS 250 J/K, T= 294 K. гхn гхn Express your answer using th ree significant figures. spontaneous or nonspontaneous

ΔΗ 133 kJ, AS 250 J/K, T-294 K. rxn rxn Express your answer using three significant figures. spontaneous or nonspontaneous

Part D ΔΗ. 133 kJ, ASo 250 J/K, T= 548 K. rxn rxn Express your answer using two significant figures. spontaneous or nonspontaneous
Answer #1

Answer A Saniu = Asynt AS surr reaction will be spontaneous if increases means As univ so entropy of unwere Now for surroundiParts As univ = -250 I -(-133(031) 294k abmw = =256 +(-_-issuetan) asumwe = 2025] Asimi 20 -spenteneous A Sunine = 2 A Sunan